The solubility of a gaseous substance A(molecular weight 26) in water is given by Henry’s law: pA=105xA ., pA in mm Hg. Convert the equilibrium relation to the following forms, a) yA=m.xA ., if the total pressure is 10 bar; (b) pA=m’CA ,CA in gmol/litre. Also write down the equilibrium relation using the mole ratio unit.


1
Expert's answer
2021-08-16T02:19:15-0400

Part a

We know that "y_A= \frac{p_A}{P} \implies p_A=y_AP"

Putting in the equation

"y_A*7500=105 x_A\\
y_A=0.014 x_A"


Part b

"p_A=105x_A\\
x_A=\frac{n_A}{n_A+n_B}\\
p_A=105 \space\frac{n_A}{n_A+n_B}\\
p_A= 105 \space\frac{c_A}{c_A+c_B}"


Part C

Given equilibrium relation

"y_A=0.014x_A"

Convert this to mole ratio unit

Y and X - mole ratio in vapour

"X= \frac{x_A}{1-x_A}\\
Y= \frac{y_A}{1-y_A}\\
X= \frac{x_A}{1+x_A}\\
Y= \frac{y_A}{1+y_A}\\
\frac{Y_A}{1+Y_A}=0.014 \frac{X_A}{1+X_A}\\"